PLease help me solve; If the width of a rectangle is 1 inch less than its length and the diagonal is 1 inch longer than its length, find the length and the width. I know w=l-1 and D=l+1 but I dont know how to fill in the blanks to solve the problem
Answer by checkley77(12844)

SUGGESTION:USE THE CAPITAL L SO YOU WON'T GET CONFUSED WITH THE SMALL l & THE 1.
W=L-1
D=L+1
L^2+W^2=D^2
L^2+(L-1)^2=(L+1)^2
L^2+L^2-2L+1=L^2+2L+1
2L^2-L^2-2L-2L+1-1=0
L^2-4L=0
L(L-4)=0
L-4=0
L=4 ANS. FOR THE LENGTH.
W=4-1=3 ANS. FOR THE WIDTH.
D=4+1=5 ANS. FOR THE DIAGONAL.
PROOF:
4^2+3^2=5^2
16+9=25
25=25